A small metal sphere weighing 10.0N is hung from a vertical spring which comes to rest after stretching 2.0cm. Determine the spring constant
1
Expert's answer
2020-03-10T11:17:52-0400

For the equilibrium:


"W=kx"

"k=\frac{W}{x}=\frac{10}{0.02}=500\frac{N}{m}"